# Friction terms in multi-fluid description of heavy-ion collisions

### v1

@@ Abstract
In multi-fluid description of heavy-ion collisions, the primary scatterings and particle production are described in terms of interaction between fluids, so called friction. These friction terms can be derived from kinetic theory, but they are not unique. We compare different approaches to derive the friction terms, introduce a new “charge transfer” friction, which allows to move charge to the midrapidity fireball, and implement them in the MUFFIN model. The charge transfer friction is more consistent with the assumption of three fluids clearly separated in momentum space, and allows better comparisons of the experimental data and underlying equation of state. It also leaves room for entropy generation due to dissipation in individual fluids, and we present the first results obtained using viscous multi-fluid dynamics.
@@ I Introduction
Heavy ion collisions allow to probe the properties of strongly interacting matter by comparing experimental measurements of the final state particle distributions to theoretical simulations of the collision dynamics. At high collision energies like those at LHC, the theoretical modeling in a multi-stage simulation framework (see e.g. Putschke:2019yrg ; Nijs:2020roc ) has had enormous success due to the fact that the physics at play allows to clearly separate different timescales of the collision: the initial fusion of gluons in the nuclei followed by a gluon dominated phase, then a non-equilibrium evolution of parton-like quasiparticles which quickly approach equilibrium such that hydrodynamics becomes applicable and finally particlization and hadronic transport. In this framework, the role of hydrodynamics is most well-established Song:2017wtw ; Florkowski:2017olj ; Heinz:2024jwu since most of the observables related to the collective behaviour are built up during the hydrodynamic stage.
At lower energies, such a separation of timescales does not appear. Since the Lorentz $\gamma$ factor is not large, the colliding nuclei have a significant longitudinal size, and the timescale of their interpenetration is comparable to that of the medium evolution. Furthermore, the processes of energy deposition, equilibration and hydrodynamic evolution all happen simultaneously. The new matter constantly coming in from the colliding nuclei causes the system to stay off equilibrium. Additionally, baryons from the colliding nuclei contribute to the plasma evolution but the baryon current decelerates much less than the energy current. Thus, the distribution of particles carrying baryon charge is significantly offset from the overall particle distribution, which is an additional non-equilibrium effect. At collision energies $\sqrt{s_{\rm NN}}\sim 10-100\,$GeV of interest in this work, this baryon transparency is manifested as a double peak structure in the final net-baryon distributions. The phenomenon of baryon transparency is already apparent in proton-proton collisions, where the distribution of outgoing baryons is strongly peaked at the edges, while outgoing pions mostly appear at midrapidity.
@@
The multi-fluid model Csernai:1982zz ; Satarov:1990uc ; Mishustin:1991sp ; Ivanov:2005yw ; Ivanov:2013wha ; Ivanov:2013yqa ; Ivanov:2013yla ; Batyuk:2016qmb ; Kozhevnikova:2020bdb ; Cimerman:2023hjw aims to describe the low energy nucleus-nucleus collision system by modeling it as multiple interacting fluids with different flow velocities. The incoming nuclei are described as baryon rich “projectile” and “target” fluids, which are decelerated by their interaction. This interaction also produces additional particles in the midrapidity region of the collision, which is modeled as a third “fireball” fluid. In momentum space, this description models the system as three distinct thermal distributions that are separated in longitudinal momentum.
@@
In this study, we introduce a charge transfer friction model in which outgoing nucleons can be assigned to any of the three fluids depending on their momentum, which more closely adheres to the picture of three separate distributions in momentum space. Furthermore, the charge transfer friction allows the build-up of finite baryon density in the fireball fluid, which makes the model suitable for exploring the equation of state at finite density. We implement the charge transfer friction in the MUFFIN model Cimerman:2023hjw , and study its behaviour via the produced charged hadron and net-proton rapidity distributions. First we briefly compare it to the two prior models, then we explore the effect of varying its free model parameters on the rapidity distributions, and finally we assess how we can best describe experimental data. We notice that the model reproduces the data better if finite shear viscosity is included, and we conclude by showing the results of the first dissipative multi-fluid calculations.
@@ II The multi-fluid model
We will now state explicitly how the projectile-target friction moves takes energy, momentum and conserved charges away from the projectile in the three different models. The effect of friction on the target is obtained by interchanging the variables related to the projectile and target ($t\leftrightarrow p$). In the Csernai model, all particles that scatter will be assigned to the fireball. This makes it easy to compute friction as the integral over all conserved quantities carried by projectile and target particles undergoing scattering:
@@
The new charge transfer friction can be understood as a generalization of the IMS friction, where outgoing nucleons at small rapidities are assigned to the fireball. This choice is more faithful to the justification of the multi-fluid model having fluids separated in momentum space. Also, it is helpful for probing the equation of state, as it allows the fireball to acquire a finite net baryon density. Specifically, outgoing nucleons with rapidities below a threshold rapidity $\beta\cdot y_{\rm in}$ are assigned to the fireball fluid and nucleons at higher rapidities to the original projectile or target fluids, where $y_{\rm in}$ is the rapidity of ingoing nucleons in the CMS frame and $\beta$ is a new parameter of the model. Therefore, the treatment of nucleons in this model is interpolated between Csernai-type friction at $\beta=1$ and IMS-type friction at $\beta=0$. The former limit faithfully reproduces Csernai friction, but the latter is not fully identical to IMS friction. This is because the correlation of the outgoing nucleon’s energy with its rapidity is not known, so the charge transfer model also introduces a second parameter $\alpha$ that quantifies what fraction of the energy coming from the decrease in nucleon rapidity is assigned to the fireball. In other words, $\alpha$ controls how much energy is used to create new particles in nucleon-nucleon scattering. In the IMS case, all nucleons stay in projectile and target, so knowledge of this correlation is not needed. The energy transfer is simply fixed by the mean energy of all outgoing pions, which is experimentally known as a function of $\sqrt{s}$. The CT friction terms are given by
$R_{pt,\rm adv}=$ $-\xi_{pt}^{2}n_{p}n_{t}V_{\rm rel}^{pt}\bar{\sigma}_{R}(s_{pt})\;,$ (13)
$F_{pt,\rm adv}^{\nu}=$ $-\xi_{pt}^{2}m_{N}n_{p}n_{t}V_{\rm rel}^{pt}\Big\{\frac{1}{2}\big[(u_{p}^{\nu}-u_{t}^{\nu})\bar{\sigma}_{P}(s_{pt})$ (14)
@@
We solve the evolution equations using MUFFIN Cimerman:2023hjw , our multi-fluid solver based on the vHLLE code Karpenko:2013wva . To describe the hadronic stage and calculate final hadron distributions, we switch to the SMASH hadron cascade SMASH:2016zqf . In the calculations presented here, we use event averaged initial states. The setup of the two colliding nuclei in the initial state and details of particlization are explained in Cimerman:2023hjw . As in Cimerman:2023hjw , we use the chiral model based Equation of State Steinheimer:2010ib . To evaluate the friction terms, we approximate the nucleon density by the net-baryon density of each fluid, and as the temperature of the pions in the fireball, we use the temperature of the fireball fluid. Likewise, the temperature required for the unification procedure (see App. C) is the temperature of the projectile and target fluids.
Figure 1: Pseudo-rapidity distributions of all charged hadrons (left) and rapidity distribution of net protons (right) with the Csernai, IMS and CT friction models using parameter values $\xi_{\rm pt}=1.0$, $\xi_{\rm f}=0.1$ for all models and $\alpha=0.7$, $\beta=0.1$ for CT friction. The data are by the PHOBOS PHOBOS:2010eyu and the NA49 NA49:1998gaz collaborations.
@@
With the inclusion of the shear stress, we need to model the shear viscosity $\eta$. We employ a recently introduced parametrisation from Jahan:2024wpj :
$\tilde{\eta}(\mu_{B})=\\ \begin{cases}\eta_{0}+(\eta_{2}-\eta_{0})\frac{\mu_{B}\text{[GeV]}}{0.2},&0<\mu_{B}\leq 0.2\text{\,GeV}\\ \eta_{2}+(\eta_{4}-\eta_{2})\frac{(\mu_{B}\text{[GeV]}-0.2)}{0.2},&0.2<\mu_{B}<0.4\text{\,GeV}\\ \eta_{4},&\mu_{B}\geq 0.4\text{\,GeV}\end{cases}$ (18)
where $\tilde{\eta}=\eta\cdot T/(\varepsilon+p)$ is the ratio of shear viscosity to enthalpy density, which is set to depend only on the baryon chemical potential. We employ the parameter values $\eta_{0}=0.045$, $\eta_{2}=0.28$ and $\eta_{4}=0.287$ obtained in the Bayesian analysis of Jahan:2024wpj .
@@ III Friction model benchmarks
Figure 2: Pseudo-rapidity distributions of all charged hadrons (left) and rapidity distribution of net protons (right) with the CT friction model. Scans over the $\beta$ and $\alpha$ parameters are conducted using $\xi_{\rm pt}=1.0$, $\xi_{\rm f}=0.1$, $\alpha=0.7$ (for top row) and $\beta=0.1$ (for bottom row). The data are by the PHOBOS PHOBOS:2010eyu and the NA49 NA49:1998gaz collaborations.
Since the different friction models lead to different longitudinal dynamics, we describe their effects on two final-state hadronic observables: pseudo-rapidity distribution of all charged hadrons and rapidity distribution of net protons. The former is sensitive to the entropy production and deposition throughout the collision, whereas the latter characterises baryon stopping. We chose the 0-5% centrality class of collisions at an energy of $\sqrt{s_{\rm NN}}=19.6\,$GeV to benchmark the friction models. At this value of the collision energy the double peak structure due to baryon transparency is already visible in the net-proton distribution. Another advantage of this choice of collision energy is that both the pseudo-rapidity distribution of charged hadrons (from the PHOBOS experiment at 19.619.6\,GeV PHOBOS:2010eyu ) and the rapidity distribution of net protons (from the NA49 experiment at 17.317.3\,GeV NA49:1998gaz ) are available at very close collision energies.
The results from different friction models are shown in Fig. 1. The Csernai friction model produces a reasonable shape of $dN_{\rm ch}/d\eta$, but fails to reproduce the experimental distribution of net protons. Because of its basic assumption that the energy, momentum and charges of all particles undergoing scatterings are immediately assigned to the fireball fluid, the transport of charges behaves exactly like the transport of energy. Therefore, there is strong baryon stopping at all collision energies. IMS-type friction improves on the prediction of the net proton distribution as it keeps all of baryon charge in the target and projectile fluids, which are only slowly decelerated by the friction. This allows to produce a double peak in the distribution as seen in Fig. 1.
@@
The charge transfer friction introduces two new parameters, $\alpha$ and $\beta$. The $\beta$ parameter sets a boundary in relative rapidity for the outgoing nucleons from the projectile-target scattering to go into either the original projectile/target fluids or to the fireball fluid. As shown in the top row of Fig. 2, with increasing $\beta$ the pseudo-rapidity distribution of all charged hadrons becomes somewhat narrower and develops a heightened plateau at mid-rapidity, because more baryons are being assigned to the fireball fluid. Correspondingly, the double peaks move farther apart and a third central peak appears in the rapidity distribution of net protons and becomes dominant at large $\beta$, which is not consistent with the experimental data at $\sqrt{s_{\rm NN}}=19.6\,$GeV.
@@
The $\alpha$ parameter controls how much energy is transferred to the fireball in nucleon-nucleon scattering, and how much of it is kept to heat up projectile and target. As shown in the bottow row of Fig. 2, with increasing $\alpha$, the charged hadron distribution becomes flatter in pseudo-rapidity, whereas the mid-rapidity dip in the distribution of net protons becomes more pronounced. The larger the value of $\alpha$, the more energy is moved to the fireball due to projectile-target friction, which also increases the entropy of the fireball. Therefore, the larger the $\alpha$, the larger the multiplicity at midrapidity. The $\alpha$ parameter also regulates the loss of energy of the projectile and target fluids relative to the loss of their momenta. At a larger value of $\alpha$, more energy is lost compared to momentum. This leads to weaker deceleration of baryon-rich fluids, which results in a more pronounced double peak in the net proton distribution. On the other hand, at very low $\alpha$, projectile and target are slowed down so much that unification is triggered for a significant portion of these baryon rich fluids and thus a large fraction of baryons is transferred to the fireball, leading to a third peak.
@@
Besides the new parameters $\alpha$ and $\beta$ of the charge transfer friction model, all models contain the parameters $\xi_{\rm pt}$ and $\xi_{\rm f}$ controlling the overall strength of the projectile-target and fireball frictions. The top row of Fig. 3 shows the results of a scan of the $\xi_{pt}$-parameter, which has strong effects on both distributions. For low values of $\xi_{\rm pt}$, little energy and momentum is transferred to the fireball, resulting in a small charged hadron multiplicity far undershooting the experimental data and slightly broader pseudo-rapidity distribution than experimentally observed. As projectile and target experience only little deceleration, the two peaks in the net-proton distribution are far apart in rapidity, close to the ingoing beam rapidities. In contrast, with increasing $\xi_{\rm pt}$, a hotter fireball results in significantly higher charged hadron production, while the two net-proton peaks become wider and move closer in rapidity eventually merging to a single peak. Choosing $\xi_{\rm pt}=2.0$ yields the best data agreement for the charged hadron distribution, but completely eliminates the double peak structure in the net-proton distribution.
@@
Collecting observations from all parameter scans, one can spot a general tension between the reproduction of experimental $dN_{\rm ch}/d\eta$ and $dN_{p-\bar{p}}/dy$. In our test simulations we have observed that typically a variation of an input parameter that improves agreement with the midrapidity $dN_{\rm ch}/d\eta$ from the data also makes the distribution of net protons narrower or eliminates the double peak structure that was measured by NA49 altogether. Therefore, in simulations of central gold-gold collisions at $\sqrt{s_{\rm NN}}=19.6\,$GeV, parameter combinations that resulted in a reasonable reproduction of experimental $dN_{\rm ch}/d\eta$ led to a shape of $dN_{p-\bar{p}}/dy$ that would be incompatible with the data. As shown in Fig. 1, even our optimal parameter combination $\alpha=0.7$, $\beta=0.1$, $\xi_{\rm pt}=1.0$, $\xi_{\rm f}=0.1$, fails to reproduce the charged hadron multiplicity at midrapidity. However, it turned out that this “failure” is not a bug, but a feature.

### v2

# Friction terms in the multi-fluid description of heavy-ion collisions

@@ Abstract
In the multi-fluid description of heavy-ion collisions, the primary scatterings and particle production are described in terms of interaction between fluids, so called friction. These friction terms can be derived from kinetic theory, but they are not unique. We compare different approaches to derive the friction terms, introduce a new “charge transfer” friction, which allows to move charge to the midrapidity fireball, and implement them in the MUFFIN model. The charge transfer friction is more consistent with the assumption of three fluids clearly separated in momentum space, and allows better comparisons of the experimental data and underlying equation of state. It also leaves room for entropy generation due to dissipation in individual fluids, and we present the first results obtained using viscous multi-fluid dynamics.
@@ I Introduction
Heavy ion collisions allow to probe the properties of strongly interacting matter by comparing experimental measurements of the final state particle distributions to theoretical simulations of the collision dynamics. At high collision energies like those at LHC, the theoretical modeling in a multi-stage simulation framework (see, e.g., Refs. Putschke:2019yrg ; Nijs:2020roc ) has had enormous success due to the fact that the physics at play allows to clearly separate different timescales of the collision: the initial fusion of gluons in the nuclei followed by a gluon dominated phase, then a non-equilibrium evolution of parton-like quasiparticles which quickly approach equilibrium such that hydrodynamics becomes applicable and finally particlization and hadronic transport. In this framework, the role of hydrodynamics is most well-established Song:2017wtw ; Florkowski:2017olj ; Heinz:2024jwu since most of the observables related to the collective behaviour are built up during the hydrodynamic stage.
At lower energies, such a separation of timescales does not appear. Since the Lorentz $\gamma$ factor is not large, the colliding nuclei have a significant longitudinal size, and the timescale of their interpenetration is comparable to that of the medium evolution. Furthermore, the processes of energy deposition, equilibration and hydrodynamic evolution all happen simultaneously. The new matter constantly coming in from the colliding nuclei causes the system to stay off equilibrium. Additionally, baryons from the colliding nuclei contribute to the plasma evolution but the baryon current decelerates much less than the energy current. Thus, the distribution of particles carrying baryon charge is significantly offset from the overall particle distribution, which is an additional non-equilibrium effect. At collision energies $\sqrt{s_{\rm NN}}\sim 10-100\,$GeV of interest in this work, this baryon transparency is manifested as a double peak structure in the final net baryon distributions. The phenomenon of baryon transparency is already apparent in proton-proton collisions, where the distribution of outgoing baryons is strongly peaked at the edges, while outgoing pions mostly appear at midrapidity.
@@
The multi-fluid model Csernai:1982zz ; Satarov:1990uc ; Mishustin:1991sp ; Ivanov:2005yw ; Ivanov:2013wha ; Ivanov:2013yqa ; Ivanov:2013yla ; Batyuk:2016qmb ; Kozhevnikova:2020bdb ; Cimerman:2023hjw ; Huovinen:2025lct aims to describe the low energy nucleus-nucleus collision system by modeling it as multiple interacting fluids with different flow velocities. The incoming nuclei are described as baryon rich “projectile” and “target” fluids, which are decelerated by their interaction. This interaction also produces additional particles in the midrapidity region of the collision, which is modeled as a third “fireball” fluid. In momentum space, this description models the system as three distinct thermal distributions that are separated in longitudinal momentum.
@@
In this study, we introduce a charge transfer friction model based on the ideas we outlined in Ref. Huovinen:2025lct . In this model, outgoing nucleons can be assigned to any of the three fluids depending on their momentum, which more closely adheres to the picture of three separate distributions in momentum space. Furthermore, the charge transfer friction allows the build-up of finite baryon density in the fireball fluid, which makes the model suitable for exploring the equation of state at finite density. We implement the charge transfer friction in the MUFFIN model Cimerman:2023hjw , and study its behaviour via the produced charged hadron and net-proton rapidity distributions. First we briefly compare it to the two prior models, then we explore the effect of varying its free model parameters on the rapidity distributions, and finally we assess how we can best describe experimental data. We notice that the model reproduces the data better if finite shear viscosity is included, and we conclude by showing the results of the first dissipative multi-fluid calculations.
@@ II The multi-fluid model
We will now state explicitly how the projectile-target friction takes energy, momentum and conserved charges away from the projectile in the three different models. For now we consider only one charge—baryon charge—but the generalization of the theory to multiple charges is straightforward. As discussed later, all fluids obey the same Equation of State with fixed $n_{Q}/n_{B}$ ratio, and therefore baryon transfer implies electric charge transfer as well. The effect of friction on the target is obtained by interchanging the variables related to the projectile and target ($t\leftrightarrow p$). In the Csernai model, all particles that scatter will be assigned to the fireball. This makes it easy to compute friction as the integral over all conserved quantities carried by projectile and target particles undergoing scattering:
@@
The new charge transfer friction can be understood as a generalization of the IMS friction, where outgoing nucleons at small rapidities are assigned to the fireball. This choice is more faithful to the justification of the multi-fluid model having fluids separated in momentum space. Also, it is helpful for probing the equation of state, as it allows the fireball to acquire a finite net-baryon density. Specifically, outgoing nucleons with rapidities below a threshold rapidity $\beta\cdot y_{\rm in}$ are assigned to the fireball fluid and nucleons at higher rapidities to the original projectile or target fluids, where $y_{\rm in}$ is the rapidity of ingoing nucleons in the CMS frame and $\beta$ is a new parameter of the model. Therefore, the treatment of nucleons in this model is interpolated between Csernai-type friction at $\beta=1$ and IMS-type friction at $\beta=0$. The former limit faithfully reproduces Csernai friction, but the latter is not fully identical to IMS friction. This is because the correlation of the outgoing nucleon’s energy with its rapidity is not known, so the charge transfer model also introduces a second parameter $\alpha$ that quantifies what fraction of the energy coming from the decrease in nucleon rapidity is assigned to the fireball. In other words, $\alpha$ controls how much energy is used to create new particles in nucleon-nucleon scattering. In the IMS case, all nucleons stay in projectile and target, so knowledge of this correlation is not needed. The energy transfer is simply fixed by the mean energy of all outgoing pions, which is experimentally known as a function of $\sqrt{s}$. The CT friction terms are given by
$R_{pt,\rm CT}=$ $-\xi_{pt}^{2}n_{p}n_{t}V_{\rm rel}^{pt}\bar{\sigma}_{R}(s_{pt})\;,$ (13)
$F_{pt,\rm CT}^{\nu}=$ $-\xi_{pt}^{2}m_{N}n_{p}n_{t}V_{\rm rel}^{pt}\Big\{\frac{1}{2}\big[(u_{p}^{\nu}-u_{t}^{\nu})\bar{\sigma}_{P}(s_{pt})$ (14)
@@
We solve the evolution equations using MUFFIN Cimerman:2023hjw , our multi-fluid solver based on the vHLLE code Karpenko:2013wva . To describe the hadronic stage and calculate final hadron distributions, we switch to the SMASH hadron cascade SMASH:2016zqf . In the calculations presented here, we use event averaged initial states. The setup of the two colliding nuclei in the initial state and details of particlization are explained in Ref. Cimerman:2023hjw . As in Ref. Cimerman:2023hjw , we use the chiral model based Equation of State (EoS) Steinheimer:2010ib constrained to zero strangeness density, $n_{S}=0$, and to electric charge density $n_{Q}=0.4\,n_{B}$. Thus the EoS depends only on energy and net-baryon densities. At particlization we employ the same constraints. To evaluate the friction terms, we approximate the nucleon density by the net-baryon density of each fluid, and as the temperature of the pions in the fireball, we use the temperature of the fireball fluid. Likewise, the temperature required for the unification procedure (see App. C) is the temperature of the projectile and target fluids.
Figure 1: Pseudorapidity distributions of all charged hadrons (left) and rapidity distribution of net protons (right) with the Csernai, IMS and CT friction models using parameter values $\xi_{\rm pt}=1.0$, $\xi_{\rm f}=0.1$ for all models and $\alpha=0.7$, $\beta=0.1$ for CT friction. The data are by the PHOBOS PHOBOS:2010eyu and the NA49 NA49:1998gaz collaborations.
@@
With the inclusion of the shear stress, we need to model the shear viscosity $\eta$. We employ a recently introduced parametrisation from Ref. Jahan:2024wpj :
$\tilde{\eta}(\mu_{B})=\\ \begin{cases}\eta_{0}+(\eta_{2}-\eta_{0})\frac{\mu_{B}\text{[GeV]}}{0.2},&0<\mu_{B}\leq 0.2\text{\,GeV},\\ \eta_{2}+(\eta_{4}-\eta_{2})\frac{(\mu_{B}\text{[GeV]}-0.2)}{0.2},&0.2<\mu_{B}<0.4\text{\,GeV},\\ \eta_{4},&\mu_{B}\geq 0.4\text{\,GeV},\end{cases}$ (18)
where $\tilde{\eta}=\eta\cdot T/(\varepsilon+p)$ is the ratio of shear viscosity to enthalpy density, which is set to depend only on the baryon chemical potential. We employ the parameter values $\eta_{0}=0.045$, $\eta_{2}=0.28$ and $\eta_{4}=0.287$ obtained in the Bayesian analysis of Jahan:2024wpj . Note that in this paper, we use the same symbol $\eta$ for both shear viscosity and pseudorapidity. To avoid confusion, we will spell out the name of the respective quantity when it appears.
@@ III Friction model benchmarks
Figure 2: Pseudorapidity distributions of all charged hadrons (left) and rapidity distribution of net protons (right) with the CT friction model. Scans over the $\beta$ and $\alpha$ parameters are conducted using $\xi_{\rm pt}=1.0$, $\xi_{\rm f}=0.1$, $\alpha=0.7$ (for top row) and $\beta=0.1$ (for bottom row). The data are by the PHOBOS PHOBOS:2010eyu and the NA49 NA49:1998gaz collaborations.
Since the different friction models lead to different longitudinal dynamics, we describe their effects on two final-state hadronic observables: pseudorapidity distribution of all charged hadrons and rapidity distribution of net protons. The former is sensitive to the entropy production and deposition throughout the collision, whereas the latter characterises baryon stopping. We chose the 0-5% centrality class of collisions at an energy of $\sqrt{s_{\rm NN}}=19.6\,$GeV to benchmark the friction models. At this value of the collision energy the double peak structure due to baryon transparency is already visible in the net-proton distribution. Another advantage of this choice of collision energy is that both the pseudorapidity distribution of charged hadrons (from the PHOBOS experiment at 19.619.6\,GeV PHOBOS:2010eyu ) and the rapidity distribution of net protons (from the NA49 experiment at 17.317.3\,GeV NA49:1998gaz ) are available at very close collision energies.
The results from different friction models are shown in Fig. 1. The Csernai friction model produces a reasonable shape of the charged hadron pseudorapidity distribution $dN_{\rm ch}/d\eta$, but fails to reproduce the experimental rapidity distribution of net protons. Because of its basic assumption that the energy, momentum and charges of all particles undergoing scatterings are immediately assigned to the fireball fluid, the transport of charges behaves exactly like the transport of energy. Therefore, there is strong baryon stopping at all collision energies. IMS-type friction improves on the prediction of the net-proton distribution as it keeps all of baryon charge in the target and projectile fluids, which are only slowly decelerated by the friction. This allows to produce a double peak in the distribution as seen in Fig. 1.
@@
The charge transfer friction introduces two new parameters, $\alpha$ and $\beta$. The $\beta$ parameter sets a boundary in relative rapidity for the outgoing nucleons from the projectile-target scattering to go into either the original projectile/target fluids or to the fireball fluid. As shown in the top row of Fig. 2, with increasing $\beta$ the pseudorapidity distribution of all charged hadrons becomes somewhat narrower and develops a heightened plateau at mid-rapidity, because more baryons are being assigned to the fireball fluid. Correspondingly, the double peaks move farther apart and a third central peak appears in the rapidity distribution of net protons and becomes dominant at large $\beta$, which is not consistent with the experimental data at $\sqrt{s_{\rm NN}}=19.6\,$GeV.
@@
The $\alpha$ parameter controls how much energy is transferred to the fireball in nucleon-nucleon scattering, and how much of it is kept to heat up projectile and target. As shown in the bottom row of Fig. 2, with increasing $\alpha$, the charged hadron distribution becomes flatter in pseudorapidity, whereas the mid-rapidity dip in the distribution of net protons becomes more pronounced. The larger the value of $\alpha$, the more energy is moved to the fireball due to projectile-target friction, which also increases the entropy of the fireball. Therefore, the larger the $\alpha$, the larger the multiplicity at midrapidity. The $\alpha$ parameter also regulates the loss of energy of the projectile and target fluids relative to the loss of their momenta. At a larger value of $\alpha$, more energy is lost compared to momentum. This leads to weaker deceleration of baryon-rich fluids, which results in a more pronounced double peak in the net-proton distribution. On the other hand, at very low $\alpha$, projectile and target are slowed down so much that unification is triggered for a significant portion of these baryon rich fluids and thus a large fraction of baryons is transferred to the fireball, leading to a third peak.
@@
Besides the new parameters $\alpha$ and $\beta$ of the charge transfer friction model, all models contain the parameters $\xi_{\rm pt}$ and $\xi_{\rm f}$ controlling the overall strength of the projectile-target and fireball frictions. The top row of Fig. 3 shows the results of a scan of the $\xi_{pt}$-parameter, which has strong effects on both distributions. For low values of $\xi_{\rm pt}$, little energy and momentum is transferred to the fireball, resulting in a small charged hadron multiplicity far undershooting the experimental data and slightly broader pseudorapidity distribution than experimentally observed. As projectile and target experience only little deceleration, the two peaks in the net-proton distribution are far apart in rapidity, close to the ingoing beam rapidities. In contrast, with increasing $\xi_{\rm pt}$, a hotter fireball results in significantly higher charged hadron production, while the two net-proton peaks become wider and move closer in rapidity eventually merging to a single peak. Choosing $\xi_{\rm pt}=2.0$ yields the best data agreement for the charged hadron distribution, but completely eliminates the double peak structure in the net-proton distribution.
@@
Collecting observations from all parameter scans, one can spot a general tension between the reproduction of the experimental charged hadron pseudorapidity distribution $dN_{\rm ch}/d\eta$ and net-proton rapidity distribution $dN_{p-\bar{p}}/dy$. In our test simulations we have observed that typically a variation of an input parameter that improves agreement with the midrapidity $dN_{\rm ch}/d\eta$ from the data also makes the distribution of net protons narrower or eliminates the double peak structure that was measured by NA49 altogether. Therefore, in simulations of central gold-gold collisions at $\sqrt{s_{\rm NN}}=19.6\,$GeV, parameter combinations that resulted in a reasonable reproduction of experimental $dN_{\rm ch}/d\eta$ led to a shape of $dN_{p-\bar{p}}/dy$ that would be incompatible with the data. As shown in Fig. 1, even our optimal parameter combination $\alpha=0.7$, $\beta=0.1$, $\xi_{\rm pt}=1.0$, $\xi_{\rm f}=0.1$, fails to reproduce the charged hadron multiplicity at midrapidity. However, it turned out that this “failure” is not a bug, but a feature.